# Enhancing Zero-Shot Facial Expression Recognition by LLM Knowledge Transfer

### v1

@@ Abstract
Current facial expression recognition (FER) models are often designed in a supervised learning manner thus are constrained by the lack of large-scale facial expression images with high-quality annotations. Consequently, these models often fail to generalize well, performing poorly on unseen images in training. Vision-language-based zero-shot models demonstrate a promising potential for addressing such challenges. However, these models lack task-specific knowledge therefore are not optimized for the nuances of recognizing facial expressions. To bridge this gap, this work proposes a novel method, Exp-CLIP, to enhance zero-shot FER by transferring the task knowledge from large language models (LLMs). Specifically, based on the pre-trained vision-language encoders, we incorporate a projection head designed to map the initial joint vision-language space into a space that captures representations of facial actions. To train this projection head for subsequent zero-shot predictions, we propose to align the projected visual representations with task-specific semantic meanings derived from the LLM encoder, and the text instruction-based strategy is employed to customize the LLM knowledge. Given unlabelled facial data and efficient training of the projection head, Exp-CLIP achieves superior zero-shot results to the CLIP models and several other large vision-language models (LVLMs) on seven in-the-wild FER datasets. The code and pre-trained models are available at https://github.com/zengqunzhao/Exp-CLIP.
@@ I Introduction
Facial expressions are a crucial aspect of human communication, allowing individuals to convey a wide range of emotions and intentions non-verbally. Facial expression recognition (FER) refers to the identification and analysis of human facial expressions to understand the underlying emotions or sentiments, which has various applications such as enabling people who are blind or have low vision to better understand the emotions of people around them [1], monitoring patient emotions for better mental health assessment [2] and improving user experience in applications and devices [3]. FER can be categorized into two types based on input distinctions: Static FER (SFER) and Dynamic FER (DFER). SFER involves analyzing individual images or snapshots of faces while DFER focuses on analyzing facial expressions across a sequence of images or frames.
@@ II-B Large Vision-Language Models
The current landscape of LVLMs can be broadly categorized into two main types. The first type involves aligning large collections of images and raw texts by utilizing two distinct encoders [20, 43, 44, 21, 24, 27, 45]. The second type integrates the visual modality into LLMs to form comprehensive LVLMs [26, 23, 25]. In the context of the first type, the pioneer work CLIP [20] maps data of different modalities, text and images, into a shared embedding space. This shared space exhibits the capability to generalize to various image classification tasks with zero- and few-shot learning. Subsequently, VideoCLIP [43] and XCLIP [44] were introduced for zero-shot video-text understanding. FaRL [21] focuses on learning universal facial representations from face-image-text pairs, while EmotionCLIP [27] is designed to extract visual emotion representations from both verbal and nonverbal communication. The recent FLAVA [24] learned a foundational language and vision representation that enables unimodal vision and language understanding as well as multimodal reasoning, all within a single pre-trained model. Regarding the LLMs-based models, Flamingo [26] bridges the powerful pre-trained vision-only and language-only models, and BLIP2 [23] further improves the capacity of the LVLMs by bridging the modality gap with a lightweight Querying Transformer and adopting more powerful LLM. The recent model LLaVa [25] was trained by fine-tuning LLMs on GPT-generated multimodal instruction-following data, demonstrating impressive multimodel chat abilities. A recent work [46] also leverages the common sense knowledge and reasoning ability of the LLaVa and shows that this knowledge can be transferred effectively to the emotion recognition problem. Nevertheless, the zero-shot FER performance of these models is challenged by substantial engineering efforts on prompting [29] and the hallucination issue inherited from LLMs decoders[30].
@@ IV-B Implementational Details
Experiment Settings: In the contrastive learning phase, the training is conducted on CAER-S [51] training set with 44,996 facial images. The parameters are optimized by the SGD optimizer with a mini-batch size of 512. The projection head in our method refers to a matrix with size $512\times 4096$ when utilising ViT-B-32/B-16 or size $768\times 4096$ when utilising ViT-L-14. The learning rate for the learnable projection head is set as $1\times 10^{-3}$. The learnable projection head is trained for 5 epochs in an end-to-end manner. Our models are trained on the Tesla A100 GPU based on the open-source PyTorch platform. We employ the Flan-T5-XXL [52] with 11B parameters as our LLM. During the zero-shot prediction, for the SFER, all the images are resized to $224\times 224$ as an input. For the DEER, all the fixed 16 frames of sequence in our experiments follow the sampling strategy in [7, 38, 12], and then resized to $224\times 224$ as inputs.
@@ IV-C Comparisons with the State-of-the-Art
We first compared our method with several large-language models in a zero-shot way, then compared with some supervised SFER and DFER state of the arts. Our results are obtained by using the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.” and the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
TABLE I: Zero-shot FER results in comparison with other visual language models. B32, B16, L14 denote ViT-B-32, ViT-B-16, and ViT-L-14, respectively. Every method uses the best-result prompt out of the three zero-shot prompts. EmoCLIP was pre-trained on MAFW [49] with facial expression clips and action unite text descriptions without categorical labels, then conducted zero-shot prediction on DFER datasets. Bold donates the best, Underline denotes the second best.
@@
Ours (B32) 36.08/42.05 31.12/31.12 27.56/27.56 39.27/36.72 25.38/29.53 22.58/24.89 17.85/21.05 31.99/33.86
Ours (B16) 49.57/55.02 38.92/38.93 33.48/33.48 39.71/53.55 25.93/31.40 25.14/28.34 18.85/23.60 31.09/32.28
Ours (L14) 57.91/65.22 44.53/44.53 38.36/38.36 46.94/55.72 40.73/47.06 28.18/31.66 24.21/27.40 40.45/41.73
TABLE II: Zero-shot FER results in comparison with CLIP model. All models use prompt “an expression of [class].” for zero-shot prediction.
@@
ViT-B-32
CLIP (plain) 39.88/35.04 32.84/32.84 27.81/27.81 37.47/27.19 23.38/19.45 21.37/17.15 18.24/18.81 30.96/29.66
CLIP (up-sampled) 34.55/36.86 28.55/28.55 25.56/25.56 31.07/24.93 18.38/18.59 16.60/19.32 14.08/19.42 24.39/25.72
Ours 41.38/49.35 33.07/33.07 29.16/29.16 39.65/45.04 28.09/25.98 23.01/20.26 19.62/22.71 31.34/31.50
$\Uparrow$ 1.50/14.31 0.23/0.23 1.35/1.35 2.18/17.85 4.71/6.53 1.64/3.11 1.38/3.90 0.38/1.84
ViT-B-16
CLIP (plain) 52.56/49.05 38.67/38.67 34.08/34.08 45.11/39.05 31.48/26.22 24.73/21.96 20.51/23.16 35.44/34.91
CLIP (up-sampled) 44.74/38.53 33.12/33.12 29.33/29.33 33.24/27.38 22.95/21.66 20.18/21.86 15.87/20.54 28.29/29.13
Ours 56.25/58.54 39.46/39.47 34.23/34.23 40.45/51.55 31.14/30.69 26.27/27.35 21.48/25.76 36.33/35.96
$\Uparrow$ 3.69/9.49 0.81/0.81 0.15/0.15 -4.66/12.50 -0.34/4.47 1.54/5.39 0.97/2.60 0.89/1.05
ViT-L-14
CLIP (plain) 49.69/37.87 37.18/37.18 33.81/33.81 42.05/26.52 32.76/20.71 20.85/14.35 19.99/16.42 33.82/32.28
CLIP (up-sampled) 48.56/39.83 39.07/39.07 34.98/34.98 43.39/27.06 33.30/25.42 22.13/18.64 22.60/21.46 33.11/33.33
Ours 57.33/61.47 42.58/42.58 37.28/37.28 51.20/55.85 38.00/38.00 27.44/26.16 25.21/29.34 37.90/37.80
$\Uparrow$ 7.64/23.60 5.40/5.40 3.47/3.47 9.15/29.33 5.24/17.29 6.59/11.81 5.22/12.92 4.08/5.52
TABLE III: Zero-shot FER results with different prompts in comparison with CLIP model. Both single prompts and prompt ensembles are investigated. For single prompts, we employ three types, labelled as P-1, P-2, and P-3 respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations, labelled as P-4 and P-5.
@@
ViT-B-32
CLIP P-1 29.62/22.33 26.12/26.12 21.93/21.93 30.32/19.80 20.88/26.63 19.56/18.19 14.93/16.21 25.74/26.51
P-2 39.88/35.04 32.84/32.84 27.81/27.81 37.47/27.19 23.38/19.45 21.37/17.15 18.24/18.81 30.96/29.66
P-3 31.27/28.42 29.29/29.29 25.53/25.53 36.06/33.92 21.70/20.61 19.19/18.03 13.34/15.46 28.59/28.87
P-4 36.69/27.12 30.35/30.35 26.73/26.73 36.19/29.45 20.97/18.21 20.69/17.29 15.96/18.42 24.82/24.67
P-5 36.18/27.44 31.67/31.67 27.66/27.66 42.01/28.05 22.75/20.22 21.35/17.59 16.36/17.19 27.99/27.82
Ours P-1 35.41/41.40 29.12/29.12 25.61/25.61 25.54/36.98 20.61/27.10 19.89/20.38 16.08/20.10 30.78/32.28
P-2 41.38/49.35 33.07/33.07 29.16/29.16 39.65/45.04 28.09/25.98 23.01/20.26 19.62/22.71 31.34/31.50
P-3 36.08/42.05 31.12/31.12 27.56/27.56 39.27/36.72 25.38/29.53 22.58/24.89 17.85/21.05 31.99/33.86
P-4 39.52/45.44 33.55/33.55 29.31/29.31 41.16/43.07 24.85/26.89 23.09/22.40 18.77/22.90 33.04/34.65
P-5 39.41/47.59 33.81/33.81 29.86/29.86 41.06/44.09 25.46/26.98 23.77/23.74 18.77/23.86 32.37/34.12
ViT-B-16
CLIP P-1 35.04/30.93 30.44/30.44 26.18/26.18 31.96/27.16 23.88/30.24 19.18/20.90 17.53/22.56 27.85/30.71
P-2 52.56/49.05 38.67/38.67 34.08/34.08 45.11/39.05 31.48/26.22 24.73/21.96 20.51/23.16 35.44/34.91
P-3 38.66/37.16 34.27/34.27 29.35/29.36 34.39/45.14 22.84/28.13 19.63/22.16 15.44/17.26 28.21/29.13
P-4 47.15/38.23 35.67/35.67 30.78/30.78 35.42/39.56 26.61/27.03 21.72/21.14 19.12/20.65 33.26/33.60
P-5 44.77/35.85 33.84/33.84 28.63/28.63 36.65/39.02 26.39/27.84 22.21/22.23 18.87/20.70 32.54/32.81
Ours P-1 43.96/49.38 34.98/34.98 30.26/30.26 38.20/45.30 23.17/29.71 23.98/25.83 19.02/22.30 39.34/40.68
P-2 56.25/58.54 39.46/39.47 34.23/34.23 40.45/51.55 31.14/30.69 26.27/27.35 21.48/25.76 36.33/35.96
P-3 49.57/55.02 38.92/38.93 33.48/33.48 39.71/53.55 25.93/31.40 25.14/28.34 18.85/23.60 31.09/32.28
P-4 54.58/59.42 39.55/39.55 33.58/33.58 42.42/56.20 30.36/32.56 27.07/28.75 21.79/26.66 35.48/36.75
P-5 52.38/59.42 40.27/40.27 34.38/34.38 43.44/55.47 30.56/33.93 27.79/29.82 21.96/28.18 36.13/36.75
ViT-L-14
CLIP P-1 48.21/35.92 36.35/36.35 29.81/29.81 46.48/32.90 27.63/24.13 20.40/15.80 19.74/16.80 29.93/28.87
P-2 49.69/37.87 37.18/37.18 33.81/33.81 42.05/26.52 32.76/20.71 20.85/14.35 19.99/16.42 33.82/32.28
P-3 47.25/40.94 34.44/34.44 29.96/29.96 33.84/46.67 31.99/38.03 21.98/28.83 20.39/25.56 35.50/38.32
P-4 51.86/38.79 39.98/39.98 34.33/34.33 44.66/33.73 35.96/32.29 22.90/20.31 23.01/25.85 34.42/34.65
P-5 54.06/42.24 39.38/39.38 33.26/33.26 44.64/36.53 36.68/38.52 24.14/26.72 24.28/28.88 35.08/35.96
Ours P-1 50.67/53.91 37.90/37.90 33.21/33.21 48.25/50.78 26.71/31.14 22.64/24.67 20.63/23.82 32.47/33.33
P-2 57.33/61.47 42.58/42.58 37.28/37.28 51.20/55.85 38.00/38.00 27.44/26.16 25.21/29.34 37.90/37.80
P-3 57.91/65.22 44.53/44.53 38.36/38.36 46.94/55.72 40.73/47.06 28.18/31.66 24.21/27.40 40.45/41.73
P-4 57.69/62.81 42.96/42.96 37.93/37.93 50.72/54.67 39.30/41.79 28.50/28.52 25.76/29.42 37.51/38.06
P-5 58.26/63.10 43.61/43.61 38.38/38.38 51.19/55.05 40.54/45.08 28.87/31.25 26.27/31.62 39.47/40.42
TABLE IV: Zero-shot FER results of our method with different projectors. All models are based on ViT-L-14, and use the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.” and the prompt “an expression of [class].” for zero-shot prediction. 2-MLP and 3-MLP denote 2-layer and 3-layer multilayer perceptron, respectively. T.E. denotes Transformer encoder. L.M. denotes our learnable linear matrix.
Methods Static FER (UAR/WAR) Dynamic FER (UAR/WAR)
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
2-MLP 49.99/59.71 34.78/34.78 29.61/29.61 40.41/47.85 31.82/34.21 24.63/24.61 18.99/20.94 31.89/33.07
3-MLP 47.12/59.97 35.44/35.44 30.88/30.88 41.58/50.49 28.36/31.52 24.51/25.16 19.45/22.81 29.79/30.18
T.E. 39.56/52.28 30.98/30.98 25.11/25.11 33.41/44.25 30.40/36.18 26.54/31.08 19.11/20.80 28.72/28.87
L.M. 57.33/61.47 42.58/42.58 37.28/37.28 51.20/55.85 38.00/38.00 27.44/26.16 25.21/29.34 37.90/37.80
@@
BLIP2 [23] is a state-of-the-art model on various vision-language tasks, to test its zero-shot performance on the FER task, inspired by work [22], we take the facial images as input and prompt BLIP2 by using: “Question: Play the role of a facial expression classification expert and select one category with the highest similarity to the input image. Here are the optional categories: ’contempt, ’fear’, ’disgust’, ’surprise’, ’anger’, ’neutral’, ’sadness’, ’happiness’. Answer:”. We then calculate the WAR and UAR based on the output. As shown in Tab. I, our results outperform the BLIP2 model on all the SFER test sets. Specifically, our ViT-B-16-based results can beat BLIP2 in zero-shot SFER. Furthermore, we also compare our results with several other visual-language models. VideoCLIP [43], XCLIP [44], and EmotionCLIP [27] are proposed for video analysis, video recognition, and video emotion recognition, respectively, in which VideoCLIP was pre-trained on videos and text transcriptions, XCLIP was pre-trained on action recognition dataset with action labels, and EmotionCLIP was pre-trained on the videos and conversations. These models all learn spatial-temporal features during the training. However, our method with the bare static model can still outperform them on four DFER test sets. Moreover, we also compared our method with CLIP betterment: FaRL and FLAVA, in which the FaRL model was pre-trained on LAION-FACE specializing in the face analysis task and FLAVA improves CLIP zero-shot performance by introducing both unimodal and multimodal architecture. The results in Tab. I demonstrate our results are superior to these two methods on all of the FER test sets. Particularly, our method outperforms EmoCLIP [28] on DFEW, FERV39k, and AFEW test sets in five metrics without explicit text descriptions, in which EmoCLIP was trained on the MAFW clips and facial action descriptions, and this indicates the effectiveness of our method.

### v2

@@ Abstract
Current facial expression recognition (FER) models are often designed in a supervised learning manner and thus are constrained by the lack of large-scale facial expression images with high-quality annotations. Consequently, these models often fail to generalize well, performing poorly on unseen images in inference. Vision-language-based zero-shot models demonstrate a promising potential for addressing such challenges. However, these models lack task-specific knowledge and therefore are not optimized for the nuances of recognizing facial expressions. To bridge this gap, this work proposes a novel method, Exp-CLIP, to enhance zero-shot FER by transferring the task knowledge from large language models (LLMs). Specifically, based on the pre-trained vision-language encoders, we incorporate a projection head designed to map the initial joint vision-language space into a space that captures representations of facial actions. To train this projection head for subsequent zero-shot predictions, we propose to align the projected visual representations with task-specific semantic meanings derived from the LLM encoder, and the text instruction-based strategy is employed to customize the LLM knowledge. Given unlabelled facial data and efficient training of the projection head, Exp-CLIP achieves superior zero-shot results to the CLIP models and several other large vision-language models (LVLMs) on seven in-the-wild FER datasets. The code and pre-trained models are available at https://github.com/zengqunzhao/Exp-CLIP.
@@ I Introduction
Facial expressions are a crucial aspect of human communication, allowing individuals to convey a wide range of emotions and intentions non-verbally. Facial expression recognition (FER) refers to the identification and analysis of human facial expressions to understand the underlying emotions or sentiments, which has various applications such as enabling people who are blind or have low vision to understand the emotions of people around them better [1], monitoring patient emotions for better mental health assessment [2] and improving user experience in applications and devices [3]. FER can be categorized into two types based on input distinctions: Static FER (SFER) and Dynamic FER (DFER). SFER involves analyzing individual images or snapshots of faces while DFER focuses on analyzing facial expressions across a sequence of images or frames.
@@ II-B Large Vision-Language Models
The current landscape of LVLMs can be broadly categorized into two main types. The first type involves aligning large collections of images and raw texts by utilizing two distinct encoders [20, 43, 44, 21, 24, 27, 45]. The second type integrates the visual modality into LLMs to form comprehensive LVLMs [26, 23, 25]. In the context of the first type, the pioneer work CLIP [20] maps data of different modalities, text and images, into a shared embedding space. This shared space exhibits the capability to generalize to various image classification tasks with zero- and few-shot learning. Subsequently, VideoCLIP [43] and XCLIP [44] were introduced for zero-shot video-text understanding. FaRL [21] focuses on learning universal facial representations from face-image-text pairs, while EmotionCLIP [27] is designed to extract visual emotion representations from both verbal and nonverbal communication. The recent FLAVA [24] learned a foundational language and vision representation that enables unimodal vision and language understanding as well as multimodal reasoning, all within a single pre-trained model. Regarding the LLMs-based models, Flamingo [26] bridges the powerful pre-trained vision-only and language-only models, and BLIP2 [23] further improves the capacity of the LVLMs by bridging the modality gap with a lightweight Querying Transformer and adopting more powerful LLM. The recent model LLaVa [25] was trained by fine-tuning LLMs on GPT-generated multimodal instruction-following data, demonstrating impressive multimodel chat abilities. A recent work [46] also leverages the common sense knowledge and reasoning ability of the LLaVa and shows that this knowledge can be transferred to the emotion recognition problem. Nevertheless, the zero-shot FER performance of these models is challenged by substantial engineering efforts on prompting [29] and the hallucination issue inherited from LLMs decoders[30].
@@ IV-B Implementational Details
Experiment Settings: In the contrastive learning phase, the training is conducted on CAER-S [51] training set with 44,996 facial images. The parameters are optimized by the SGD optimizer with a mini-batch size of 512. The projection head in our method refers to a matrix with size $512\times 4096$ when utilising ViT-B-32/B-16 or size $768\times 4096$ when utilising ViT-L-14. The learning rate for the learnable projection head is set as $1\times 10^{-3}$. The learnable projection head is trained for 5 epochs in an end-to-end manner. Our models are trained on the Tesla A100 GPU based on the open-source PyTorch platform. We employ the Flan-T5-XXL [52] with 11B parameters as our LLM. During the zero-shot prediction, for the SFER, all the images are resized to $224\times 224$ as an input. For the DEER, all the fixed 16 frames of sequence in our experiments follow the sampling strategy in [7, 38, 12], and then resized to $224\times 224$ as inputs. To obtain more stable and reliable results, we train models three times with different random seeds and then use the average as the final result.
@@ IV-C Comparisons with the State-of-the-Art
We first compared our method with several large-language models in a zero-shot way, then compared with some supervised SFER and DFER state of the arts. Our results are obtained using the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”, and the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
TABLE I: Zero-shot FER results in comparison with other visual language models. B/32, B/16, L/14 denote ViT-B/32, ViT-B/16, and ViT-L/14, respectively. EmoCLIP was pre-trained on MAFW [49] with facial expression clips and action unit text descriptions without categorical labels. Bold donates the best, Underline denotes the second best.
@@
Exp-CLIP B/32 (Ours) 39.56/42.14 31.73/31.73 27.63/27.63 37.83/37.04 24.25/25.87 21.48/21.25 17.53/20.27 29.72/31.23
Exp-CLIP B/16 (Ours) 48.96/54.50 39.98/39.98 34.40/34.40 40.81/53.02 27.12/31.75 24.78/27.99 19.17/23.35 32.84/33.86
Exp-CLIP L/14 (Ours) 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
TABLE II: Zero-shot FER results in comparison with CLIP model. All models use the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
@@
ViT-B/32
CLIP [20] 34.97/30.05 29.29/29.29 25.53/25.53 36.05/33.89 22.08/21.30 18.12/17.38 13.39/15.48 28.59/28.87
Exp-CLIP (Ours) 39.56/42.14 31.73/31.73 27.63/27.63 37.83/37.04 24.25/25.87 21.48/21.25 17.53/20.27 29.72/31.23
$\Uparrow$ 4.59/12.09 2.44/2.44 2.10/2.10 1.78/3.15 2.17/4.57 3.36/3.87 4.14/4.79 1.13/2.36
ViT-B/16
CLIP [20] 38.66/37.16 34.35/34.35 29.38/29.38 34.33/45.14 22.92/28.33 19.81/22.51 15.45/17.25 28.21/29.13
Exp-CLIP (Ours) 48.96/54.50 39.98/39.98 34.40/34.40 40.81/53.02 27.12/31.75 24.78/27.99 19.17/23.35 32.84/33.86
$\Uparrow$ 9.11/15.75 5.63/5.63 5.02/5.02 6.48/7.88 4.20/3.42 4.97/5.48 3.72/6.10 4.63/4.73
ViT-L/14
CLIP [20] 47.22/41.13 34.46/34.47 29.96/29.96 33.82/46.67 32.02/38.12 21.97/28.99 20.42/25.58 35.50/38.32
Exp-CLIP (Ours) 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
$\Uparrow$ 11.48/24.24 9.81/9.80 8.48/8.47 14.46/8.75 8.14/8.97 5.48/2.08 3.53/1.40 3.22/2.10
TABLE III: Zero-shot FER results with different prompts in comparison with CLIP model. Both single prompts and prompt ensembles are investigated. We employ three types for single prompts, labelled as P-1, P-2, and P-3 respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations, labelled as P-4 and P-5.
@@
ViT-B/32
CLIP P-1 28.62/23.57 26.09/26.09 21.86/21.86 30.29/19.76 21.07/26.94 19.34/18.82 14.99/16.35 25.74/26.51
P-2 40.04/35.76 32.84/32.84 27.86/27.86 37.40/27.03 23.22/19.85 20.71/16.99 18.26/18.85 30.97/29.66
P-3 34.97/30.05 29.29/29.29 25.53/25.53 36.05/33.89 22.08/21.30 18.12/17.38 13.39/15.48 28.59/28.87
P-4 35.58/25.36 29.69/29.69 25.53/25.53 34.30/25.28 20.58/18.36 18.56/15.83 15.59/17.16 28.87/28.87
P-5 37.73/30.28 31.61/31.61 27.31/27.31 39.85/29.45 23.75/22.48 20.90/17.84 17.51/19.73 27.26/27.30
Ours P-1 35.88/41.59 29.89/29.90 25.93/25.93 28.81/38.38 21.15/26.73 20.03/19.76 15.77/19.84 29.60/31.32
P-2 42.12/49.12 33.61/33.61 29.31/29.31 39.82/45.55 26.38/25.90 22.53/20.16 18.99/22.14 31.98/31.93
P-3 39.56/42.14 31.73/31.73 27.63/27.63 37.83/37.04 24.25/25.87 21.48/21.25 17.53/20.27 29.72/31.23
P-4 41.04/45.47 32.56/32.56 28.43/28.43 36.24/42.93 23.57/24.00 21.19/19.98 18.14/21.74 33.69/35.35
P-5 41.13/47.63 33.62/33.63 29.29/29.29 38.38/43.89 24.48/24.85 21.98/20.46 18.40/22.97 32.91/34.65
ViT-B/16
CLIP P-1 32.78/29.27 30.41/30.41 26.26/26.26 31.98/27.16 23.87/30.44 19.23/21.13 17.56/22.61 27.85/30.71
P-2 53.37/50.52 38.61/38.61 34.03/34.03 45.10/39.02 31.48/26.45 24.80/22.26 20.58/23.22 35.44/34.91
P-3 39.85/38.75 34.35/34.35 29.38/29.38 34.33/45.14 22.92/28.33 19.81/22.51 15.45/17.25 28.21/29.13
P-4 42.69/34.71 35.83/35.84 30.48/30.48 39.25/35.64 25.71/28.85 20.19/20.31 17.60/20.52 32.56/32.81
P-5 41.74/37.42 34.49/34.50 29.70/29.71 34.42/36.95 27.89/29.38 23.44/23.37 19.98/22.24 33.12/34.12
Ours P-1 41.80/48.58 35.43/35.44 30.52/30.52 38.72/45.29 24.65/31.10 24.42/27.31 19.43/22.37 37.52/39.02
P-2 55.88/60.21 39.75/39.76 34.44/34.44 42.42/53.25 30.78/31.52 26.72/28.26 21.74/26.31 35.98/35.52
P-3 48.96/54.50 39.98/39.98 34.40/34.40 40.81/53.02 27.12/31.75 24.78/27.99 19.17/23.35 32.84/33.86
P-4 52.42/57.26 41.10/41.10 35.05/35.05 43.30/52.94 28.52/32.55 26.19/29.03 21.35/26.84 36.85/37.71
P-5 50.82/57.66 40.76/40.76 34.88/34.88 43.05/54.30 30.79/34.59 27.53/30.05 22.17/27.40 37.02/38.50
ViT-L/14
CLIP P-1 48.20/38.14 36.24/36.24 29.76/29.76 46.59/32.99 27.74/24.43 19.75/15.59 19.77/16.79 29.93/28.87
P-2 49.37/39.34 37.35/37.35 33.86/33.86 42.06/26.49 32.68/20.83 20.95/14.54 20.02/16.46 33.82/32.28
P-3 47.22/41.13 34.46/34.47 29.96/29.96 33.82/46.67 32.02/38.12 21.97/28.99 20.42/25.58 35.50/38.32
P-4 49.79/36.34 39.41/39.41 33.53/33.53 43.65/30.22 35.39/29.31 21.97/18.85 21.84/22.86 31.61/31.50
P-5 52.21/41.36 39.72/39.73 32.98/32.98 43.28/31.40 35.85/35.40 23.54/24.10 23.20/26.99 34.08/34.65
Ours P-1 49.08/53.13 38.05/38.06 32.93/32.93 48.55/49.76 26.76/31.65 22.31/25.27 20.69/23.92 32.45/33.95
P-2 56.17/60.77 42.68/42.68 37.46/37.46 51.54/55.15 39.04/38.48 26.45/25.10 24.81/28.47 38.50/38.58
P-3 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
P-4 56.15/60.23 43.48/43.48 38.03/38.03 50.97/53.30 38.15/39.75 27.14/27.57 25.14/27.72 40.17/41.29
P-5 56.52/61.41 43.68/43.68 38.11/38.11 49.93/53.37 39.02/43.57 28.70/31.70 25.94/30.99 40.31/42.00
@@
BLIP2 [23] is a state-of-the-art model on various vision-language tasks, to test its zero-shot performance on the FER task, inspired by work [22], we take the facial images as input and prompt BLIP2 by using: “Question: Play the role of a facial expression classification expert and select one category with the highest similarity to the input image. Here are the optional categories: ‘contempt’, ‘fear’, ‘disgust’, ‘surprise’, ‘anger’, ‘neutral’, ‘sadness’, ‘happiness’. Answer:”. We then calculate the WAR and UAR based on the output. As shown in Tab. I, our results outperform the BLIP2 model on all the FER test sets. Specifically, our ViT-B-16-based zero-shot results can beat BLIP2 in most SFER test sets.
Furthermore, we also compare our results with several other visual-language models. VideoCLIP [43], XCLIP [44], and EmotionCLIP [27] are proposed for video analysis, video recognition, and video emotion recognition, respectively, in which VideoCLIP was pre-trained on video clips and text transcriptions, XCLIP was pre-trained on action recognition dataset with action labels, and EmotionCLIP was pre-trained on the videos and conversations. These models all learn spatial-temporal features during the training. However, our method with the bare static model can still outperform them on four DFER test sets. Moreover, we also compared our method with CLIP betterment: FaRL and FLAVA, in which the FaRL model was pre-trained on LAION-FACE specializing in the face analysis task and FLAVA improves CLIP zero-shot performance by introducing both unimodal and multimodal architecture. The results in Tab. I demonstrate our results are superior to these two methods on all of the FER test sets. Compared with the EmoCLIP [28], which was trained on the MAFW clips and facial action descriptions, our method outperforms it on DFEW, FERV39k, and AFEW test sets in five metrics without explicit text descriptions, indicating the effectiveness of the proposed method.